OpenStudy (anonymous):

please help write the inverse of the area function of a circle and leave the answer in terms of pi. kinda confused here

OpenStudy (accessdenied):

So, do you know the area function of the circle? As in, given a radius r, the function gives the area of the circle of radius r.

OpenStudy (anonymous):

yes i know that

OpenStudy (anonymous):

a=pir^2

OpenStudy (accessdenied):

Yup. Then, to find an inverse function, we would like to solve for r in terms of A. Is that part clear as well?

OpenStudy (anonymous):

it says inverse in terms of pi though

OpenStudy (accessdenied):

What it means by that is, we will keep the pi as it is in the equation, rather than say rounding it to 3.14 or something: So if this were our equation (circumference in terms of pi and r): C = 2 pi r We solve for r by dividing by 2 pi: C / (2 pi) = r We leave pi as is and now have the function r in terms of pi and C. Does that make more sense about what the question is asking?

OpenStudy (anonymous):

ok i see

OpenStudy (accessdenied):

Awesome. Then you see how to do this problem? I can go over the work if you like.

OpenStudy (anonymous):

sqrt(a/pi)=r ?

OpenStudy (accessdenied):

Looks good! And worthy of note is that because the radius is always nonnegative (r > 0), it is the positive square root rather than negative.

OpenStudy (anonymous):

ok thanks!